\noindent

\begin{thebibliography}{999}                                                                                              %


\bibitem {IN-Penrose:1969pc}R.~Penrose, ``Gravitational collapse: The role of
general relativity,'' Riv.\ Nuovo Cim.\ \textbf{1}, 252 (1969)
[Gen.\ Rel.\ Grav.\ \textbf{34}, 1141 (2002)].

\bibitem {IN-Christodoulou:1970wf}D.~Christodoulou, ``Reversible and
irreversible transforations in black hole physics,''
Phys.\ Rev.\ Lett.\ \textbf{25}, 1596 (1970). doi:10.1103/PhysRevLett.25.1596

\bibitem {IN-Bardeen:1970zz}J.~M.~Bardeen, ``Kerr Metric Black Holes,'' Nature
\textbf{226}, 64 (1970). doi:10.1038/226064a0

\bibitem {IN-Christodoulou:1972kt}D.~Christodoulou and R.~Ruffini,
``Reversible transformations of a charged black hole,'' Phys.\ Rev.\ D
\textbf{4}, 3552 (1971). doi:10.1103/PhysRevD.4.3552

\bibitem {IN-Bekenstein:1972tm}J.~D.~Bekenstein, ``Black holes and the second
law,'' Lett.\ Nuovo Cim.\ \textbf{4}, 737 (1972). doi:10.1007/BF02757029

\bibitem {IN-Bekenstein:1973ur}J.~D.~Bekenstein, \textquotedblleft Black holes
and entropy,\textquotedblright\ Phys.\ Rev.\ D \textbf{7}, 2333 (1973). doi:10.1103/PhysRevD.7.2333

\bibitem {IN-Bardeen:1973gs}J.~M.~Bardeen, B.~Carter and S.~W.~Hawking,
\textquotedblleft The Four laws of black hole mechanics,\textquotedblright%
\ Commun.\ Math.\ Phys.\ \textbf{31}, 161 (1973). doi:10.1007/BF01645742

\bibitem {IN-Maldacena:1997re}J.~M.~Maldacena, \textquotedblleft The Large N
limit of superconformal field theories and supergravity,\textquotedblright%
\ Int.\ J.\ Theor.\ Phys.\ \textbf{38}, 1113 (1999)
[Adv.\ Theor.\ Math.\ Phys.\ \textbf{2}, 231 (1998)]
doi:10.1023/A:1026654312961, 10.4310/ATMP.1998.v2.n2.a1 [hep-th/9711200].

\bibitem {IN-Chamblin:1999tk}A.~Chamblin, R.~Emparan, C.~V.~Johnson and
R.~C.~Myers, \textquotedblleft Charged AdS black holes and catastrophic
holography,\textquotedblright\ Phys.\ Rev.\ D \textbf{60}, 064018 (1999)
doi:10.1103/PhysRevD.60.064018 [hep-th/9902170].

\bibitem {IN-Chamblin:1999hg}A.~Chamblin, R.~Emparan, C.~V.~Johnson and
R.~C.~Myers, \textquotedblleft Holography, thermodynamics and fluctuations of
charged AdS black holes,\textquotedblright\ Phys.\ Rev.\ D \textbf{60}, 104026
(1999) doi:10.1103/PhysRevD.60.104026 [hep-th/9904197].

\bibitem {IN-Caldarelli:1999xj}M.~M.~Caldarelli, G.~Cognola and D.~Klemm,
``Thermodynamics of Kerr-Newman-AdS black holes and conformal field
theories,'' Class.\ Quant.\ Grav.\ \textbf{17}, 399 (2000)
doi:10.1088/0264-9381/17/2/310 [hep-th/9908022].

\bibitem {IN-Cai:2001dz}R.~G.~Cai, \textquotedblleft Gauss-Bonnet black holes
in AdS spaces,\textquotedblright\ Phys.\ Rev.\ D \textbf{65}, 084014 (2002)
doi:10.1103/PhysRevD.65.084014 [hep-th/0109133].

\bibitem {IN-Kubiznak:2012wp}D.~Kubiznak and R.~B.~Mann, \textquotedblleft P-V
criticality of charged AdS black holes,\textquotedblright\ JHEP \textbf{1207},
033 (2012) doi:10.1007/JHEP07(2012)033 [arXiv:1205.0559 [hep-th]].

\bibitem {IN-Gunasekaran:2012dq}S.~Gunasekaran, R.~B.~Mann and D.~Kubiznak,
\textquotedblleft Extended phase space thermodynamics for charged and rotating
black holes and Born-Infeld vacuum polarization,\textquotedblright\ JHEP
\textbf{1211}, 110 (2012) doi:10.1007/JHEP11(2012)110 [arXiv:1208.6251 [hep-th]].

\bibitem {IN-Wei:2012ui}S.~W.~Wei and Y.~X.~Liu, \textquotedblleft Critical
phenomena and thermodynamic geometry of charged Gauss-Bonnet AdS black
holes,\textquotedblright\ Phys.\ Rev.\ D \textbf{87}, no. 4, 044014 (2013)
doi:10.1103/PhysRevD.87.044014 [arXiv:1209.1707 [gr-qc]].

\bibitem {IN-Cai:2013qga}R.~G.~Cai, L.~M.~Cao, L.~Li and R.~Q.~Yang, ``P-V
criticality in the extended phase space of Gauss-Bonnet black holes in AdS
space,'' JHEP \textbf{1309}, 005 (2013) doi:10.1007/JHEP09(2013)005
[arXiv:1306.6233 [gr-qc]].

\bibitem {IN-Xu:2014kwa}W.~Xu and L.~Zhao, ``Critical phenomena of static
charged AdS black holes in conformal gravity,'' Phys.\ Lett.\ B \textbf{736},
214 (2014) doi:10.1016/j.physletb.2014.07.019 [arXiv:1405.7665 [gr-qc]].

\bibitem {IN-Frassino:2014pha}A.~M.~Frassino, D.~Kubiznak, R.~B.~Mann and
F.~Simovic, ``Multiple Reentrant Phase Transitions and Triple Points in
Lovelock Thermodynamics,'' JHEP \textbf{1409}, 080 (2014)
doi:10.1007/JHEP09(2014)080 [arXiv:1406.7015 [hep-th]].

\bibitem {IN-Dehghani:2014caa}M.~H.~Dehghani, S.~Kamrani and A.~Sheykhi,
``$P-V$ criticality of charged dilatonic black holes,'' Phys.\ Rev.\ D
\textbf{90}, no. 10, 104020 (2014) doi:10.1103/PhysRevD.90.104020
[arXiv:1505.02386 [hep-th]].

\bibitem {IN-Hennigar:2015esa}R.~A.~Hennigar, W.~G.~Brenna and R.~B.~Mann,
\textquotedblleft$P-v$ criticality in quasitopological
gravity,\textquotedblright\ JHEP \textbf{1507}, 077 (2015)
doi:10.1007/JHEP07(2015)077 [arXiv:1505.05517 [hep-th]].

\bibitem {IN-Wang:2018xdz}P.~Wang, H.~Wu and H.~Yang, ``Thermodynamics and
Phase Transitions of Nonlinear Electrodynamics Black Holes in an Extended
Phase Space,'' JCAP \textbf{1904}, 052 (2019)
doi:10.1088/1475-7516/2019/04/052 [arXiv:1808.04506 [gr-qc]].

\bibitem {IN-Wu:2019xse}D.~Wu, P.~Wu, H.~Yu and S.~Q.~Wu, ``Notes on the
thermodynamics of superentropic AdS black holes,'' Phys.\ Rev.\ D
\textbf{101}, no. 2, 024057 (2020) doi:10.1103/PhysRevD.101.024057

\bibitem {IN-Wald}R. Wald, \textquotedblleft Gedanken experiments to destroy a
black hole,\textquotedblright\ Ann. Phys.\textbf{ 82}, 548 (1974) doi:10.1016/0003-4916(74)90125-0

\bibitem {IN-Hubeny:1998ga}V.~E.~Hubeny, ``Overcharging a black hole and
cosmic censorship,'' Phys.\ Rev.\ D \textbf{59}, 064013 (1999)
doi:10.1103/PhysRevD.59.064013 [gr-qc/9808043].

\bibitem {IN-Jacobson:2009kt}T.~Jacobson and T.~P.~Sotiriou, ``Over-spinning a
black hole with a test body,'' Phys.\ Rev.\ Lett.\ \textbf{103}, 141101 (2009)
Erratum: [Phys.\ Rev.\ Lett.\ \textbf{103}, 209903 (2009)]
doi:10.1103/PhysRevLett.103.209903, 10.1103/PhysRevLett.103.141101
[arXiv:0907.4146 [gr-qc]].

\bibitem {IN-Saa:2011wq}A.~Saa and R.~Santarelli, \textquotedblleft Destroying
a extremal Kerr-Newman black hole,\textquotedblright\ Phys.\ Rev.\ D
\textbf{84}, 027501 (2011) doi:10.1103/PhysRevD.84.027501 [arXiv:1105.3950 [gr-qc]].

\bibitem {IN-Hod:2008zza}S.~Hod, ``Weak Cosmic Censorship: As Strong as
Ever,'' Phys.\ Rev.\ Lett.\ \textbf{100}, 121101 (2008)
doi:10.1103/PhysRevLett.100.121101 [arXiv:0805.3873 [gr-qc]].

\bibitem {IN-Barausse:2010ka}E.~Barausse, V.~Cardoso and G.~Khanna, ``Test
bodies and naked singularities: Is the self-force the cosmic censor?,''
Phys.\ Rev.\ Lett.\ \textbf{105}, 261102 (2010)
doi:10.1103/PhysRevLett.105.261102 [arXiv:1008.5159 [gr-qc]].

\bibitem {IN-Barausse:2011vx}E.~Barausse, V.~Cardoso and G.~Khanna, ``Testing
the Cosmic Censorship Conjecture with point particles: the effect of radiation
reaction and the self-force,'' Phys.\ Rev.\ D \textbf{84}, 104006 (2011)
doi:10.1103/PhysRevD.84.104006 [arXiv:1106.1692 [gr-qc]].

\bibitem {IN-Zimmerman:2012zu}P.~Zimmerman, I.~Vega, E.~Poisson and R.~Haas,
``Self-force as a cosmic censor,'' Phys.\ Rev.\ D \textbf{87}, no. 4, 041501
(2013) doi:10.1103/PhysRevD.87.041501 [arXiv:1211.3889 [gr-qc]].

\bibitem {IN-Colleoni:2015afa}M.~Colleoni and L.~Barack, ``Overspinning a Kerr
black hole: the effect of self-force,'' Phys.\ Rev.\ D \textbf{91}, 104024
(2015) doi:10.1103/PhysRevD.91.104024 [arXiv:1501.07330 [gr-qc]].

\bibitem {IN-Colleoni:2015ena}M.~Colleoni, L.~Barack, A.~G.~Shah and M.~van de
Meent, ``Self-force as a cosmic censor in the Kerr overspinning problem,''
Phys.\ Rev.\ D \textbf{92}, no. 8, 084044 (2015)
doi:10.1103/PhysRevD.92.084044 [arXiv:1508.04031 [gr-qc]].

\bibitem {IN-Matsas:2007bj}G.~E.~A.~Matsas and A.~R.~R.~da Silva,
``Overspinning a nearly extreme charged black hole via a quantum tunneling
process,'' Phys.\ Rev.\ Lett.\ \textbf{99}, 181301 (2007)
doi:10.1103/PhysRevLett.99.181301 [arXiv:0706.3198 [gr-qc]].

\bibitem {IN-Richartz:2008xm}M.~Richartz and A.~Saa, ``Overspinning a nearly
extreme black hole and the Weak Cosmic Censorship conjecture,'' Phys.\ Rev.\ D
\textbf{78}, 081503 (2008) doi:10.1103/PhysRevD.78.081503 [arXiv:0804.3921 [gr-qc]].

\bibitem {IN-Isoyama:2011ea}S.~Isoyama, N.~Sago and T.~Tanaka, ``Cosmic
censorship in overcharging a Reissner-Nordstr\'{o}m black hole via charged
particle absorption,'' Phys.\ Rev.\ D \textbf{84}, 124024 (2011)
doi:10.1103/PhysRevD.84.124024 [arXiv:1108.6207 [gr-qc]].

\bibitem {IN-Gao:2012ca}S.~Gao and Y.~Zhang, ``Destroying extremal Kerr-Newman
black holes with test particles,'' Phys.\ Rev.\ D \textbf{87}, no. 4, 044028
(2013) doi:10.1103/PhysRevD.87.044028 [arXiv:1211.2631 [gr-qc]].

\bibitem {IN-Hod:2013vj}S.~Hod, ``Cosmic Censorship: Formation of a Shielding
Horizon Around a Fragile Horizon,'' Phys.\ Rev.\ D \textbf{87}, no. 2, 024037
(2013) doi:10.1103/PhysRevD.87.024037 [arXiv:1302.6658 [gr-qc]].

\bibitem {IN-Duztas:2013wua}K.~Duztas and I.~Semiz, ``Cosmic Censorship, Black
Holes and Integer-spin Test Fields,'' Phys.\ Rev.\ D \textbf{88}, no. 6,
064043 (2013) doi:10.1103/PhysRevD.88.064043 [arXiv:1307.1481 [gr-qc]].

\bibitem {IN-Siahaan:2015ljs}H.~M.~Siahaan, ``Destroying Kerr-Sen black
holes,'' Phys.\ Rev.\ D \textbf{93}, no. 6, 064028 (2016)
doi:10.1103/PhysRevD.93.064028 [arXiv:1512.01654 [gr-qc]].

\bibitem {IN-Natario:2016bay}J.~Natario, L.~Queimada and R.~Vicente, ``Test
fields cannot destroy extremal black holes,''
Class.\ Quant.\ Grav.\ \textbf{33}, no. 17, 175002 (2016)
doi:10.1088/0264-9381/33/17/175002 [arXiv:1601.06809 [gr-qc]].

\bibitem {IN-Duztas:2016xfg}K.~Duztas, ``Overspinning BTZ black holes with
test particles and fields,'' Phys.\ Rev.\ D \textbf{94}, no. 12, 124031 (2016)
doi:10.1103/PhysRevD.94.124031 [arXiv:1701.07241 [gr-qc]].

\bibitem {IN-Revelar:2017sem}K.~S.~Revelar and I.~Vega, ``Overcharging
higher-dimensional black holes with point particles,'' Phys.\ Rev.\ D
\textbf{96}, no. 6, 064010 (2017) doi:10.1103/PhysRevD.96.064010
[arXiv:1706.07190 [gr-qc]].

\bibitem {IN-Sorce:2017dst}J.~Sorce and R.~M.~Wald, ``Gedanken experiments to
destroy a black hole. II. Kerr-Newman black holes cannot be overcharged or
overspun,'' Phys.\ Rev.\ D \textbf{96}, no. 10, 104014 (2017)
doi:10.1103/PhysRevD.96.104014 [arXiv:1707.05862 [gr-qc]].

\bibitem {IN-Husain:2017cmj}V.~Husain and S.~Singh, ``Penrose inequality in
anti--de Sitter space,'' Phys.\ Rev.\ D \textbf{96}, no. 10, 104055 (2017)
doi:10.1103/PhysRevD.96.104055 [arXiv:1709.02395 [gr-qc]].

\bibitem {IN-Crisford:2017gsb}T.~Crisford, G.~T.~Horowitz and J.~E.~Santos,
``Testing the Weak Gravity - Cosmic Censorship Connection,'' Phys.\ Rev.\ D
\textbf{97}, no. 6, 066005 (2018) doi:10.1103/PhysRevD.97.066005
[arXiv:1709.07880 [hep-th]].

\bibitem {IN-Gwak:2017kkt}B.~Gwak, ``Thermodynamics with Pressure and Volume
under Charged Particle Absorption,'' JHEP \textbf{1711}, 129 (2017)
doi:10.1007/JHEP11(2017)129 [arXiv:1709.08665 [gr-qc]].

\bibitem {IN-An:2017phb}J.~An, J.~Shan, H.~Zhang and S.~Zhao,
``Five-dimensional Myers-Perry black holes cannot be overspun in gedanken
experiments,'' Phys.\ Rev.\ D \textbf{97}, no. 10, 104007 (2018)
doi:10.1103/PhysRevD.97.104007 [arXiv:1711.04310 [hep-th]].

\bibitem {IN-Ge:2017vun}B.~Ge, Y.~Mo, S.~Zhao and J.~Zheng,
``Higher-dimensional charged black holes cannot be over-charged by gedanken
experiments,'' Phys.\ Lett.\ B \textbf{783}, 440 (2018)
doi:10.1016/j.physletb.2018.07.015 [arXiv:1712.07342 [hep-th]].

\bibitem {IN-Yu:2018eqq}T.~Y.~Yu and W.~Y.~Wen, ``Cosmic censorship and Weak
Gravity Conjecture in the Einstein--Maxwell-dilaton theory,'' Phys.\ Lett.\ B
\textbf{781}, 713 (2018) doi:10.1016/j.physletb.2018.04.060 [arXiv:1803.07916 [gr-qc]].

\bibitem {IN-Gwak:2018akg}B.~Gwak, ``Weak Cosmic Censorship Conjecture in
Kerr-(Anti-)de Sitter Black Hole with Scalar Field,'' JHEP \textbf{1809}, 081
(2018) doi:10.1007/JHEP09(2018)081 [arXiv:1807.10630 [gr-qc]].

\bibitem {IN-Gim:2018axz}Y.~Gim and B.~Gwak, ``Charged black hole in gravity's
rainbow: Violation of weak cosmic censorship,'' Phys.\ Lett.\ B \textbf{794},
122 (2019) doi:10.1016/j.physletb.2019.05.039 [arXiv:1808.05943 [gr-qc]].

\bibitem {IN-Chen:2018yah}D.~Chen, ``Weak cosmic censorship conjecture in BTZ
black holes with scalar fields,'' Chin.\ Phys.\ C \textbf{44}, no. 1, 015101
(2020) doi:10.1088/1674-1137/44/1/015101 [arXiv:1812.03459 [gr-qc]].

\bibitem{IN-Zeng:2019jta}
X.~X.~Zeng and H.~Q.~Zhang,
``Thermodynamics and weak cosmic censorship conjecture in Born-Infeld-anti-de Sitter black holes,''
Chin. Phys. C \textbf{45}, no.2, 025112 (2021)
doi:10.1088/1674-1137/abd088
[arXiv:1901.04247 [hep-th]].

\bibitem {IN-Chen:2019nsr}D.~Chen, W.~Yang and X.~Zeng, ``Thermodynamics and
weak cosmic censorship conjecture in Reissner-Nordstr$\ddot{o}$m anti-de
Sitter black holes with scalar field,'' Nucl.\ Phys.\ B \textbf{946}, 114722
(2019) doi:10.1016/j.nuclphysb.2019.114722 [arXiv:1901.05140 [hep-th]].

\bibitem {IN-Gwak:2019asi}B.~Gwak, ``Weak Cosmic Censorship with Pressure and
Volume in Charged Anti-de Sitter Black Hole under Charged Scalar Field,'' JCAP
\textbf{1908}, 016 (2019) doi:10.1088/1475-7516/2019/08/016 [arXiv:1901.05589 [gr-qc]].

\bibitem {IN-Zeng:2019jrh}X.~X.~Zeng, Y.~W.~Han and D.~Y.~Chen,
``Thermodynamics and weak cosmic censorship conjecture of BTZ black holes in
extended phase space,'' Chin.\ Phys.\ C \textbf{43}, no. 10, 105104 (2019)
doi:10.1088/1674-1137/43/10/105104 [arXiv:1901.08915 [gr-qc]].

\bibitem {IN-Chen:2019pdj}D.~Chen, ``Thermodynamics and weak cosmic censorship
conjecture in extended phase spaces of anti-de Sitter black holes with
particles' absorption,'' Eur.\ Phys.\ J.\ C \textbf{79}, no. 4, 353 (2019)
doi:10.1140/epjc/s10052-019-6874-5 [arXiv:1902.06489 [hep-th]].

\bibitem {IN-Wang:2019jzz}P.~Wang, H.~Wu and H.~Yang, ``Thermodynamics of
nonlinear electrodynamics black holes and the validity of weak cosmic
censorship at charged particle absorption,'' Eur.\ Phys.\ J.\ C \textbf{79},
no. 7, 572 (2019). doi:10.1140/epjc/s10052-019-7090-z
%%CITATION = doi:10.1140/epjc/s10052-019-7090-z;%%


\bibitem {IN-Zeng:2019hux}X.~X.~Zeng, X.~Y.~Hu and K.~J.~He, ``Weak cosmic
censorship conjecture with pressure and volume in the Gauss-Bonnet AdS black
hole,'' Nucl.\ Phys.\ B \textbf{949}, 114823 (2019)
doi:10.1016/j.nuclphysb.2019.114823 [arXiv:1905.07750 [hep-th]].

\bibitem {IN-Hong:2019yiz}W.~Hong, B.~Mu and J.~Tao, ``Thermodynamics and weak
cosmic censorship conjecture in the charged RN-AdS black hole surrounded by
quintessence under the scalar field,'' Nucl.\ Phys.\ B \textbf{949}, 114826
(2019) doi:10.1016/j.nuclphysb.2019.114826 [arXiv:1905.07747 [gr-qc]].

\bibitem {IN-Hu:2019rpw}T.~T.~Hu, Y.~Song, S.~Sun, H.~B.~Li and Y.~Q.~Wang,
``Weak cosmic censorship in Born-Infeld electrodynamics and bound on
charge-to-mass ratio,'' Eur.\ Phys.\ J.\ C \textbf{80}, no. 2, 147 (2020)
doi:10.1140/epjc/s10052-020-7703-6 [arXiv:1906.00235 [hep-th]].

\bibitem {IN-Hu:2019qzw}S.~Q.~Hu, Y.~C.~Ong and D.~N.~Page, ``No evidence for
violation of the second law in extended black hole thermodynamics,''
Phys.\ Rev.\ D \textbf{100}, no. 10, 104022 (2019)
doi:10.1103/PhysRevD.100.104022 [arXiv:1906.05870 [gr-qc]].

\bibitem {IN-Mu:2019bim}B.~Mu and J.~Tao, ``Minimal Length Effect on
Thermodynamics and Weak Cosmic Censorship Conjecture in anti-de Sitter Black
Holes via Charged Particle Absorption,'' arXiv:1906.10544 [gr-qc].

\bibitem {IN-He:2019fti}K.~J.~He, X.~Y.~Hu and X.~X.~Zeng, ``Weak cosmic
censorship conjecture and thermodynamics in quintessence AdS black hole under
charged particle absorption,'' Chin.\ Phys.\ C \textbf{43}, no. 12, 125101
(2019) doi:10.1088/1674-1137/43/12/125101 [arXiv:1906.10531 [gr-qc]].

\bibitem {IN-Gan:2019jac}Q.~Gan, G.~Guo, P.~Wang and H.~Wu, ``Strong cosmic
censorship for a scalar field in a Born-Infeld--de Sitter black hole,''
Phys.\ Rev.\ D \textbf{100}, no. 12, 124009 (2019)
doi:10.1103/PhysRevD.100.124009 [arXiv:1907.04466 [hep-th]].

\bibitem {IN-Zeng:2019baw}X.~X.~Zeng and X.~Y.~Hu, ``Thermodynamics and weak
cosmic censorship conjecture with pressure in the rotating BTZ black holes,''
arXiv:1908.03845 [gr-qc].


\bibitem{IN-He:2019kws}
K.~J.~He, G.~P.~Li and X.~Y.~Hu,
``Violations of the weak cosmic censorship conjecture in the higher dimensional $f(R)$ black holes with pressure,''
Eur. Phys. J. C \textbf{80}, no.3, 209 (2020)
doi:10.1140/epjc/s10052-020-7669-4
[arXiv:1909.09956 [hep-th]].

\bibitem{IN-Hu:2019zxr}
S.~Q.~Hu, B.~Liu, X.~M.~Kuang and R.~H.~Yue,
``Revisiting black hole thermodynamics in massive gravity: charged particle absorption and shell of dust falling,''
Chin. Phys. C \textbf{44}, 105107 (2020)
doi:10.1088/1674-1137/abab8a
[arXiv:1910.04437 [gr-qc]].

\bibitem{IN-Gwak:2019rcz}
B.~Gwak,
``Weak Cosmic Censorship in Kerr-Sen Black Hole under Charged Scalar Field,''
JCAP \textbf{03}, 058 (2020)
doi:10.1088/1475-7516/2020/03/058
[arXiv:1910.13329 [gr-qc]].

\bibitem{IN-Gan:2019ibg}
Q.~Gan, P.~Wang, H.~Wu and H.~Yang,
``Strong Cosmic Censorship for a Scalar Field in an Einstein-Maxwell-Gauss-Bonnet-de Sitter Black Hole,''
Chin. Phys. C \textbf{45}, no.2, 025103 (2021)
doi:10.1088/1674-1137/abccaf
[arXiv:1911.10996 [gr-qc]].


\bibitem{IN-Hong:2020zcf}
W.~Hong, B.~Mu and J.~Tao,
``Testing the weak cosmic censorship conjecture in torus-like black hole under charged scalar field,''
Int. J. Mod. Phys. D \textbf{29}, no.12, 2050078 (2020)
doi:10.1142/S0218271820500789
[arXiv:2001.09008 [physics.gen-ph]].

\bibitem{IN-Gwak:2020zht}
B.~Gwak,
``Thermodynamics in Rotating Anti-de Sitter Black Holes with Massive Scalar Field in Three Dimensions,''
Chin. Phys. C \textbf{44}, no.12, 12 (2020)
doi:10.1088/1674-1137/abb656
[arXiv:2002.09729 [gr-qc]].

\bibitem {IN-Zhang:2013tba}Y.~Zhang and S.~Gao, \textquotedblleft Testing
cosmic censorship conjecture near extremal black holes with cosmological
constants,\textquotedblright\ Int.\ J.\ Mod.\ Phys.\ D \textbf{23}, 1450044
(2014) doi:10.1142/S0218271814500448 [arXiv:1309.2027 [gr-qc]].

\bibitem {IN-York:1986it}J.~W.~York, Jr., ``Black hole thermodynamics and the
Euclidean Einstein action,'' Phys.\ Rev.\ D \textbf{33}, 2092 (1986). doi:10.1103/PhysRevD.33.2092

\bibitem {IN-Braden:1990hw}H.~W.~Braden, J.~D.~Brown, B.~F.~Whiting and
J.~W.~York, Jr., ``Charged black hole in a grand canonical ensemble,''
Phys.\ Rev.\ D \textbf{42}, 3376 (1990). doi:10.1103/PhysRevD.42.3376

\bibitem {IN-Carlip:2003ne}S.~Carlip and S.~Vaidya, ``Phase transitions and
critical behavior for charged black holes,''
Class.\ Quant.\ Grav.\ \textbf{20}, 3827 (2003)
doi:10.1088/0264-9381/20/16/319 [gr-qc/0306054].

\bibitem {IN-Lundgren:2006kt}A.~P.~Lundgren, ``Charged black hole in a
canonical ensemble,'' Phys.\ Rev.\ D \textbf{77}, 044014 (2008)
doi:10.1103/PhysRevD.77.044014 [gr-qc/0612119].

\bibitem {IN-Lu:2010xt}J.~X.~Lu, S.~Roy and Z.~Xiao, ``Phase transitions and
critical behavior of black branes in canonical ensemble,'' JHEP \textbf{1101},
133 (2011) doi:10.1007/JHEP01(2011)133 [arXiv:1010.2068 [hep-th]].

\bibitem {IN-Wu:2011yu}C.~Wu, Z.~Xiao and J.~Xu, ``Bubbles and Black Branes in
Grand Canonical Ensemble,'' Phys.\ Rev.\ D \textbf{85}, 044009 (2012)
doi:10.1103/PhysRevD.85.044009 [arXiv:1108.1347 [hep-th]].

\bibitem {IN-Lu:2012rm}J.~X.~Lu, R.~Wei and J.~Xu, ``The phase structure of
black D1/D5 (F/NS5) system in canonical ensemble,'' JHEP \textbf{1212}, 012
(2012) doi:10.1007/JHEP12(2012)012 [arXiv:1210.0708 [hep-th]].

\bibitem {IN-Lu:2013nt}J.~X.~Lu and R.~Wei, ``Modulating the phase structure
of black D6 branes in canonical ensemble,'' JHEP \textbf{1304}, 100 (2013)
doi:10.1007/JHEP04(2013)100 [arXiv:1301.1780 [hep-th]].

\bibitem {IN-Zhou:2015yxa}D.~Zhou and Z.~Xiao, \textquotedblleft Phase
structures of the black D$p$-D($p+4$)-brane system in various ensembles I:
thermal stability,\textquotedblright\ JHEP \textbf{1507}, 134 (2015)
doi:10.1007/JHEP07(2015)134 [arXiv:1502.00261 [hep-th]].

\bibitem {IN-Xiao:2015bha}Z.~Xiao and D.~Zhou, \textquotedblleft Phase
structures of the black D$p$-D($p+4$)-brane system in various ensembles II:
electrical and thermodynamic stability,\textquotedblright\ JHEP \textbf{1509},
028 (2015) doi:10.1007/JHEP09(2015)028 [arXiv:1507.02088 [hep-th]].

\bibitem{IN-Wang:2019urm}
P.~Wang, H.~Yang and S.~Ying,
``Thermodynamics and phase transition of a Gauss-Bonnet black hole in a cavity,''
Phys. Rev. D \textbf{101}, no.6, 064045 (2020)
doi:10.1103/PhysRevD.101.064045
[arXiv:1909.01275 [gr-qc]].

\bibitem {IN-Sanchis-Gual:2015lje}N.~Sanchis-Gual, J.~C.~Degollado,
P.~J.~Montero, J.~A.~Font and C.~Herdeiro, ``Explosion and Final State of an
Unstable Reissner-Nordstrom Black Hole,'' Phys.\ Rev.\ Lett.\ \textbf{116},
no. 14, 141101 (2016) doi:10.1103/PhysRevLett.116.141101 [arXiv:1512.05358 [gr-qc]].

\bibitem {IN-Sanchis-Gual:2016tcm}N.~Sanchis-Gual, J.~C.~Degollado,
C.~Herdeiro, J.~A.~Font and P.~J.~Montero, ``Dynamical formation of a
Reissner-Nordstrom black hole with scalar hair in a cavity,'' Phys.\ Rev.\ D
\textbf{94}, no. 4, 044061 (2016) doi:10.1103/PhysRevD.94.044061
[arXiv:1607.06304 [gr-qc]].

\bibitem {IN-Basu:2016srp}P.~Basu, C.~Krishnan and P.~N.~B.~Subramanian,
``Hairy Black Holes in a Box,'' JHEP \textbf{1611}, 041 (2016)
doi:10.1007/JHEP11(2016)041 [arXiv:1609.01208 [hep-th]].

\bibitem {IN-Peng:2017gss}Y.~Peng, B.~Wang and Y.~Liu, ``On the thermodynamics
of the black hole and hairy black hole transitions in the asymptotically flat
spacetime with a box,'' Eur.\ Phys.\ J.\ C \textbf{78}, no. 3, 176 (2018)
doi:10.1140/epjc/s10052-018-5652-0 [arXiv:1708.01411 [hep-th]].

\bibitem {IN-Peng:2017squ}Y.~Peng, ``Studies of a general flat space/boson
star transition model in a box through a language similar to holographic
superconductors,'' JHEP \textbf{1707}, 042 (2017) doi:10.1007/JHEP07(2017)042
[arXiv:1705.08694 [hep-th]].

\bibitem {IN-Peng:2018abh}Y.~Peng, ``Scalar field configurations supported by
charged compact reflecting stars in a curved spacetime,'' Phys.\ Lett.\ B
\textbf{780}, 144 (2018) doi:10.1016/j.physletb.2018.02.068 [arXiv:1801.02495 [gr-qc]].

\bibitem {IN-Peng:2019qrl}Y.~Peng, ``Scalarization of compact stars in the
scalar-Gauss-Bonnet gravity,'' JHEP \textbf{1912}, 064 (2019)
doi:10.1007/JHEP12(2019)064 [arXiv:1910.13718 [gr-qc]].


\bibitem{IN-Peng:2020zmu}
Y.~Peng,
``Analytical investigations on formations of hairy neutral reflecting shells in the scalar-Gauss\textendash{}Bonnet gravity,''
Eur. Phys. J. C \textbf{80}, no.3, 202 (2020)
doi:10.1140/epjc/s10052-020-7778-0
[arXiv:2002.01892 [gr-qc]].

\bibitem {IN-Wang:2019kxp}P.~Wang, H.~Wu and H.~Yang, ``Thermodynamics and
Phase Transition of a Nonlinear Electrodynamics Black Hole in a Cavity,'' JHEP
\textbf{1907}, 002 (2019) doi:10.1007/JHEP07(2019)002 [arXiv:1901.06216 [gr-qc]].


\bibitem{IN-Liang:2019dni}
K.~Liang, P.~Wang, H.~Wu and M.~Yang,
``Phase structures and transitions of Born\textendash{}Infeld black holes in a grand canonical ensemble,''
Eur. Phys. J. C \textbf{80}, no.3, 187 (2020)
doi:10.1140/epjc/s10052-020-7750-z
[arXiv:1907.00799 [gr-qc]].

\bibitem{IN-Wang:2019cax}
P.~Wang, H.~Wu and H.~Yang,
``Thermodynamic Geometry of AdS Black Holes and Black Holes in a Cavity,''
Eur. Phys. J. C \textbf{80}, no.3, 216 (2020)
doi:10.1140/epjc/s10052-020-7776-2
[arXiv:1910.07874 [gr-qc]].

\bibitem {IN-Simovic:2018tdy}F.~Simovic and R.~B.~Mann, \textquotedblleft
Critical Phenomena of Charged de Sitter Black Holes in
Cavities,\textquotedblright\ Class.\ Quant.\ Grav.\ \textbf{36}, no. 1, 014002
(2019) doi:10.1088/1361-6382/aaf445 [arXiv:1807.11875 [gr-qc]].


\bibitem{IN-Haroon:2020vpr}
S.~Haroon, R.~A.~Hennigar, R.~B.~Mann and F.~Simovic,
``Thermodynamics of Gauss-Bonnet-de Sitter Black Holes,''
Phys. Rev. D \textbf{101}, 084051 (2020)
doi:10.1103/PhysRevD.101.084051
[arXiv:2002.01567 [gr-qc]].


\bibitem {HJE-Benrong:2014woa}B.~Mu, P.~Wang and H.~Yang, \textquotedblleft
Covariant GUP Deformed Hamilton-Jacobi Method,\textquotedblright\ Adv.\ High
Energy Phys.\ \textbf{2017}, 3191839 (2017) doi:10.1155/2017/3191839
[arXiv:1408.5055 [gr-qc]].

\bibitem {HJE-Tao:2017mpe}J.~Tao, P.~Wang and H.~Yang, \textquotedblleft Black
hole radiation with modified dispersion relation in tunneling paradigm: Static
frame,\textquotedblright\ Nucl.\ Phys.\ B \textbf{922}, 346 (2017)
doi:10.1016/j.nuclphysb.2017.06.022 [arXiv:1505.03045 [gr-qc]].

\bibitem {Con-Jiang:2019ige}J.~Jiang, B.~Deng and Z.~Chen, ``Static charged
dilaton black hole cannot be overcharged by gedanken experiments,'' Phys. Rev.
D \textbf{100}, no.6, 066024 (2019) doi:10.1103/PhysRevD.100.066024
[arXiv:1909.02219 [hep-th]].

\bibitem {Con-Jiang:2019vww}J.~Jiang, X.~Liu and M.~Zhang, ``Examining the
weak cosmic censorship conjecture by gedanken experiments for Kerr-Sen black
holes,'' Phys. Rev. D \textbf{100}, no.8, 084059 (2019)
doi:10.1103/PhysRevD.100.084059 [arXiv:1910.04060 [hep-th]].

\bibitem {Con-Wang:2019bml}X.~Y.~Wang and J.~Jiang, ``Examining the weak
cosmic censorship conjecture of RN-AdS black holes via the new version of the
gedanken experiment,'' JCAP \textbf{07}, 052 (2020)
doi:10.1088/1475-7516/2020/07/052 [arXiv:1911.03938 [hep-th]].
\end{thebibliography}

\documentclass[aps,preprint]{revtex4}%
\usepackage{amsfonts}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{graphicx}%
\begin{document}
\preprint{CTP-SCU/2020003}
\title{Validity of Thermodynamic Laws and Weak Cosmic Censorship for AdS Black Holes
and Black Holes in a Cavity}
\author{Peng Wang$^{a}$}
\email{pengw@scu.edu.cn}
\author{Houwen Wu$^{a}$}
\email{iverwu@scu.edu.cn}
\author{Shuxuan Ying$^{a,b}$}
\email{ysxuan@cqu.edu.cn}
\affiliation{$^{a}$Center for Theoretical Physics, College of Physics, Sichuan University,
Chengdu, 610064, PR China}
\affiliation{$^{b}$Department of Physics, Chongqing University, Chongqing, 401331, China}

\begin{abstract}
By throwing a test charged particle into a Reissner-Nordstrom (RN) black hole,
we test the validity of the first and second laws of thermodynamics and weak
cosmic censorship conjecture (WCCC) with two types of boundary conditions,
i.e., the asymptotically anti-de Sitter (AdS) space and a Dirichlet cavity
wall placed in the asymptotically flat space. For the RN-AdS black hole, the
second law of thermodynamics is satisfied, and the WCCC is violated for both
extremal and near-extremal black holes. For the RN black hole in a cavity, the
entropy can either increase or decrease depending on the change in the charge,
and WCCC is satisfied/violated for the extremal/near-extremal black hole. Our
results indicate that there may be a connection between the black hole
thermodynamics and the boundary condition imposed on the black hole.

\end{abstract}
\keywords{}\maketitle
\tableofcontents

\section{Introduction}

Studying thermodynamic properties of black holes can have a deep impact upon
the understanding of quantum gravity. Penrose first noticed that a particle
can extract energy from a black hole with an ergosphere
\cite{IN-Penrose:1969pc}, which led to the discovery of the irreducible mass
\cite{IN-Christodoulou:1970wf,IN-Bardeen:1970zz,IN-Christodoulou:1972kt}. The
square of the irreducible mass of a black hole can be interpreted as the black
hole entropy \cite{IN-Bekenstein:1972tm,IN-Bekenstein:1973ur}. Later,
analogous to the laws of thermodynamics, the four laws of black hole mechanics
were proposed \cite{IN-Bardeen:1973gs}. With the advent of the AdS/CFT
correspondence \cite{IN-Maldacena:1997re}, there has been much interest in
studying thermodynamics and phase structure of AdS black holes
\cite{IN-Chamblin:1999tk,IN-Chamblin:1999hg,IN-Caldarelli:1999xj,IN-Cai:2001dz,IN-Kubiznak:2012wp,IN-Wei:2012ui,IN-Gunasekaran:2012dq,IN-Cai:2013qga,IN-Xu:2014kwa,IN-Frassino:2014pha,IN-Dehghani:2014caa,IN-Hennigar:2015esa,IN-Wang:2018xdz,IN-Wu:2019xse}%
, where some intriguing phase behavior, e.g., reentrant phase transitions and
tricritical points, was found to be present. It is also worth pointing out
that black holes can become thermally stable in AdS space since the AdS
boundary acts as a reflecting wall.

Along with the development of black hole thermodynamics, the weak cosmic
censorship conjecture (WCCC) was proposed to hide singularities by event
horizons \cite{IN-Penrose:1969pc}. So if the WCCC is valid, the singularities
cannot be seen by the observers at the future null infinity. To test the
validity of the WCCC, Wald constructed a gedanken experiment to destroy an
extremal Kerr-Newman black hole by overcharging or overspinning the black hole
via throwing a test particle into it \cite{IN-Wald}. Nevertheless, the
extremal Kerr-Newman black hole was shown to be incapable of capturing
particles with sufficient charge or angular momentum to overcharge or overspin
the black hole. Later, near-extremal charged/rotating black holes were found
to be overcharged/overspun by absorbing a particle
\cite{IN-Hubeny:1998ga,IN-Jacobson:2009kt,IN-Saa:2011wq}, and hence the WCCC
is violated. However, subsequent researches showed that the WCCC might be
still valid if the backreaction and self-force effects were considered
\cite{IN-Hod:2008zza,IN-Barausse:2010ka,IN-Barausse:2011vx,IN-Zimmerman:2012zu,IN-Colleoni:2015afa,IN-Colleoni:2015ena}%
. Since there is still a lack of the general proof of the WCCC, its validity
has been tested in various black holes
\cite{IN-Matsas:2007bj,IN-Richartz:2008xm,IN-Isoyama:2011ea,IN-Gao:2012ca,IN-Hod:2013vj,IN-Duztas:2013wua,IN-Siahaan:2015ljs,IN-Natario:2016bay,IN-Duztas:2016xfg,IN-Revelar:2017sem,IN-Sorce:2017dst,IN-Husain:2017cmj,IN-Crisford:2017gsb,IN-Gwak:2017kkt,IN-An:2017phb,IN-Ge:2017vun,IN-Yu:2018eqq,IN-Gwak:2018akg,IN-Gim:2018axz,IN-Chen:2018yah,IN-Zeng:2019jta,IN-Chen:2019nsr,IN-Gwak:2019asi,IN-Zeng:2019jrh,IN-Chen:2019pdj,IN-Wang:2019jzz,IN-Zeng:2019hux,IN-Hong:2019yiz,IN-Hu:2019rpw,IN-Hu:2019qzw,IN-Mu:2019bim,IN-He:2019fti,IN-Gan:2019jac,IN-Zeng:2019baw,IN-He:2019kws,IN-Hu:2019zxr,IN-Gwak:2019rcz,IN-Gan:2019ibg,IN-Hong:2020zcf,IN-Gwak:2020zht}%
. In particular, the thermodynamics and WCCC have been considered for a
Reissner-Nordstrom (RN)-AdS black via the charged particle absorption in the
normal and extended phase spaces \cite{IN-Gwak:2017kkt,IN-Zhang:2013tba}. In
the normal phase space, in which the cosmological constant is fixed, it showed
that the first and second laws of thermodynamics are satisfied while the WCCC
is violated even for an extremal RN-AdS black hole.

Instead of the AdS boundary, York showed that placing Schwarzschild black
holes inside a cavity, on the wall of which the metric is fixed, can make them
thermally stable \cite{IN-York:1986it}. Thermodynamics and phase structure of
RN black holes in a cavity were studied in a grand canonical ensemble
\cite{IN-Braden:1990hw} and a canonical ensemble
\cite{IN-Carlip:2003ne,IN-Lundgren:2006kt}. It was found that the
Schwarzschild and RN black holes in a cavity have quite similar phase
structure and transition to these of the AdS counterparts. Afterwards, various
black brane systems
\cite{IN-Lu:2010xt,IN-Wu:2011yu,IN-Lu:2012rm,IN-Lu:2013nt,IN-Zhou:2015yxa,IN-Xiao:2015bha}%
, a Gauss-Bonnet black hole \cite{IN-Wang:2019urm}, hairy black holes
\cite{IN-Sanchis-Gual:2015lje,IN-Sanchis-Gual:2016tcm,IN-Basu:2016srp,IN-Peng:2017gss}
and Boson stars
\cite{IN-Peng:2017squ,IN-Peng:2018abh,IN-Peng:2019qrl,IN-Peng:2020zmu} in a
cavity were extensively investigated, and it also showed that the behavior of
the gravity systems in a cavity is strikingly similar to that of the
counterparts in AdS gravity. However, we have recently studied phase structure
of Born-Infeld black holes enclosed in a cavity
\cite{IN-Wang:2019kxp,IN-Liang:2019dni} and thermodynamic geometry of RN black
holes in a cavity \cite{IN-Wang:2019cax}, and found their behavior has
dissimilarities from that of the corresponding AdS black holes. Note that
thermodynamics and critical behavior of de Sitter black holes in a cavity were
investigated in \cite{IN-Simovic:2018tdy,IN-Haroon:2020vpr}.

Although there have been a lot of work in progress on the thermodynamic laws
and WCCC for various black holes of different theories of gravity in
spacetimes with differing asymptotics, little is known about the second law of
thermodynamics and WCCC for a black hole enclosed in a cavity. Since RN-AdS
black holes and RN black holes in a cavity are thermally stable, they provide
appropriate scenarios to explore whether or not the thermodynamic laws and
WCCC are sensitive to the boundary condition of black holes. To this end, we
study the thermodynamic laws and WCCC for a RN black hole in a cavity in this paper.

The rest of this paper is organized as follows. In section \ref{Sec:RNAdSBH},
we review the discussion of the thermodynamic laws and WCCC for a RN-AdS black
hole to be self-contained and introduce the method used in this paper. The
thermodynamic laws and WCCC of a RN black hole in a cavity are then tested via
the absorption of a charged particle in section \ref{Sec:RNBHC}. We summarize
our results with a brief discussion in section \ref{Sec:Con}. For simplicity,
we set $G=\hbar=c=k_{B}=1$ in this paper.

\section{RN-AdS Black Hole}

\label{Sec:RNAdSBH}

In this section, we discuss the first and second laws of thermodynamics and
WCCC for a RN-AdS black hole by throwing a test particle into the black hole.
First, we consider the motion of a test particle of energy $E$, charge $q$ and
mass $m$ in a $4$-dimensional charged static black hole with the line
element,
\begin{equation}
ds^{2}=-f\left(  r\right)  dt^{2}+\frac{1}{f\left(  r\right)  }dr^{2}%
+r^{2}\left(  d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)  ,
\end{equation}
and the electromagnetic potential $A_{\mu}$,
\begin{equation}
A_{\mu}=A_{t}\left(  r\right)  \delta_{\mu t}.
\end{equation}
We also suppose that the outermost horizon of the black hole is at $r=r_{+}$,
where $f\left(  r_{+}\right)  =0$. In \cite{IN-Wang:2019jzz}, the
Hamilton-Jacobi equation of the test particle was given by%
\begin{equation}
-\frac{\left[  E+qA_{t}\left(  r\right)  \right]  ^{2}}{f\left(  r\right)
}+\frac{\left[  P^{r}\left(  r\right)  \right]  ^{2}}{f\left(  r\right)
}+\frac{L^{2}}{r^{2}}=m^{2}\text{,} \label{eq:HJE}%
\end{equation}
where $L$ and $P^{r}\left(  r\right)  $ are the particle's angular momentum
and radial momentum, respectively. It was shown in
\cite{HJE-Benrong:2014woa,HJE-Tao:2017mpe} that $P^{r}\left(  r_{+}\right)  $
is finite and proportional to the Hawking temperature of the black hole. Eqn.
$\left(  \ref{eq:HJE}\right)  $ gives
\begin{equation}
E=-qA_{t}\left(  r\right)  +\sqrt{f\left(  r\right)  \left(  m^{2}+\frac
{L^{2}}{r^{2}}\right)  +\left[  P^{r}\left(  r\right)  \right]  ^{2}},
\label{eq:EqA}%
\end{equation}
where we choose the positive sign in front of the square root since the energy
of the particle is required to be a positive value
\cite{IN-Christodoulou:1970wf,IN-Christodoulou:1972kt}. At the horizon
$r=r_{+}$, the above equation reduces to%
\begin{equation}
E=q\Phi+\left\vert P^{r}\left(  r_{+}\right)  \right\vert , \label{eq:EP}%
\end{equation}
where $\Phi\equiv-A_{t}\left(  r_{+}\right)  $ is the electric potential of
the black hole. Eqn. $\left(  \ref{eq:EP}\right)  $ relates the energy of the
particle to its radial momentum and potential energy just before the particle
enters the horizon.

To check whether a particle can reach or exist near the black hole
horizon, we can rewrite eqn. $\left(  \ref{eq:EqA}\right)  $ as the radial
equation of motion,%
\begin{equation}
\left(  \frac{dr}{d\tau}\right)  ^{2}=\frac{\left[  E+qA_{t}\left(  r\right)
\right]  ^{2}}{m^{2}}-f\left(  r\right)  \left(  1+\frac{L^{2}}{m^{2}r^{2}%
}\right)  ,\label{eq:drdtau}%
\end{equation}
where we use $P^{r}\left(  r\right)  =mdr/d\tau$, and $\tau$ is the affine
parameter along the worldline. Note that a particle can exist in the region
where $\left(  dr/d\tau\right)  ^{2}\geq0$, and $dr/d\tau=0$ corresponds to a
turning point. Specifically for a particle existing at the event horizon, eqn.
$\left(  \ref{eq:drdtau}\right)  $ gives that $\left(  dr/d\tau\right)
^{2}|_{r=r_{+}}\geq0$, and hence $E\geq q\Phi$. Furthermore, if the particle
falls into the black hole, one has $\left(  dr/d\tau\right)  ^{2}>0$ at the
event horizon, which leads to $E>q\Phi$. In short, for a particle of energy
$E$ and charge $q$ around a black hole of potential $\Phi$, $E>q\Phi$ provides
a lower bound $E_{\text{low}}$ on $E$, which ensures that the particle exists
near the event horizon and gets absorbed by the black hole.

For a RN-AdS black hole, the metric function $f\left(  r\right)  $ and
electric potential $\Phi$ are%
\begin{equation}
f\left(  r;M,Q\right)  =1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{r^{2}}{l^{2}%
}\text{ and }\Phi=\frac{Q}{r_{+}}, \label{eq:AdSmetric}%
\end{equation}
respectively, where $M$ and $Q$ are the mass and charge of the black hole,
respectively, and $l$ is the AdS radius. Here the parameters $M$ and $Q$ are
put explicitly as arguments of the function $f\left(  r\right)  $ for later
convenience. With fixed charge $Q$, the mass $M_{e}\left(  Q\right)  $ and
horizon radius $r_{e}\left(  Q\right)  $ of the extremal RN-AdS black hole are
determined by $f\left(  r_{e}\left(  Q\right)  ;M,Q\right)  =df\left(
r;M,Q\right)  /dr|_{r=r_{e}\left(  Q\right)  }=0$, which gives%
\begin{align}
M_{e}\left(  Q\right)   &  =\frac{\sqrt{6}l}{18}\left(  2+\sqrt{1+12Q^{2}%
/l^{2}}\right)  \sqrt{\sqrt{1+12Q^{2}/l^{2}}-1},\nonumber\\
r_{e}\left(  Q\right)   &  =\frac{l}{\sqrt{6}}\sqrt{\sqrt{1+12Q^{2}/l^{2}}-1}.
\end{align}
Like a RN black hole, when $M\geq$ $M_{e}\left(  Q\right)  $, the RN-AdS black
hole solution possesses the event horizon at $r=r_{+}\left(  M,Q\right)  $,
which is obtained by solving $f\left(  r;M,Q\right)  =0$
\cite{IN-Zhang:2013tba}. Otherwise, the event horizon disappears, and a naked
singularity emerges, which leads to the violation of the WCCC. If the event
horizon exists, one can define the black hole's temperature $T$ and entropy
$S$ as%
\begin{equation}
T=\frac{1}{4\pi}\frac{\partial f\left(  r;M,Q\right)  }{\partial r}%
|_{r=r_{+}\left(  M,Q\right)  }\text{ and }S=\pi r_{+}^{2}\left(  M,Q\right)
. \label{eq:AdSTandS}%
\end{equation}

Suppose one starts with an initial black hole of mass $M$ and charge $Q$ with
$M\geq$ $M_{e}\left(  Q\right)  $ and throws a test particle of energy $E\ll
M$ and charge $q\ll Q$ into the black hole. After the particle is absorbed,
the final configuration has mass $M+dM$ and charge $Q+dQ$. The energy and
charge conservation of the absorbing process gives%
\begin{equation}
dM=E\text{ and }dQ=q\text{,} \label{eq:AdSdMdQ}%
\end{equation}
where $E$ and $q$ are related via eqn. $\left(  \ref{eq:EP}\right)  $. If
$M+E\geq$ $M_{e}\left(  Q+q\right)  $, there exits an event horizon in the
final black hole solution, which hides the naked singularity. However for
$M+E<$ $M_{e}\left(  Q+q\right)  $, the naked singularity can be seen by
distant observers due to the absence of the event horizon.

We first check the first and second laws of thermodynamics for a RN-AdS black
hole during the absorption. In this case, the final black hole solution should
possess an event horizon at $r=r_{+}\left(  M,Q\right)  +dr_{+}$ in order that
thermodynamic variables are well defined.\ So the horizon radius, mass and
charge of the final black hole satisfy%
\begin{equation}
f\left(  r_{+}\left(  M,Q\right)  +dr_{+};M+dM,Q+dQ\right)  =0\text{,}%
\end{equation}
which leads to%
\begin{equation}
\frac{\partial f\left(  r;M,Q\right)  }{\partial r}|_{r=r_{+}\left(
M,Q\right)  }dr_{+}+\frac{\partial f\left(  r;M,Q\right)  }{\partial
M}|_{r=r_{+}\left(  M,Q\right)  }dM+\frac{\partial f\left(  r;M,Q\right)
}{\partial Q}|_{r=r_{+}\left(  M,Q\right)  }dQ=0\text{.}%
\end{equation}
Using eqns. $\left(  \ref{eq:EP}\right)  $, $\left(  \ref{eq:AdSmetric}%
\right)  $, $\left(  \ref{eq:AdSTandS}\right)  $ and $\left(  \ref{eq:AdSdMdQ}%
\right)  $, one finds that the above equation gives
\begin{equation}
\left\vert P^{r}\left(  r_{+}\left(  M,Q\right)  \right)  \right\vert
=TdS\text{.} \label{eq:AdSPT}%
\end{equation}
For an extremal black hole with $T=0$, since $P^{r}\left(  r_{+}\left(
M,Q\right)  \right)  \propto T$, eqn. $\left(  \ref{eq:AdSPT}\right)  $ is
trivial. Nevertheless, for a non-extremal RN-AdS black hole with $T>0$, the
variation of entropy is%
\begin{equation}
dS=\frac{\left\vert P^{r}\left(  r_{+}\left(  M,Q\right)  \right)  \right\vert
}{T}>0\text{,}%
\end{equation}
which means the second law of thermodynamics is satisfied. Moreover, plugging
eqn. $\left(  \ref{eq:AdSdMdQ}\right)  $ and $\left(  \ref{eq:AdSPT}\right)  $
into eqn. $\left(  \ref{eq:EP}\right)  $ yields the first law of
thermodynamics:%
\begin{equation}
dM=\Phi dQ+TdS\text{.}%
\end{equation}

To test the WCCC, we consider an extremal or near-extremal RN-AdS black hole
and check whether throwing a charged particle can overcharge the black hole.
To overcharge the black hole, the final configuration should exceeds
extremality:%
\begin{equation}
M+E<M_{e}\left(  Q+q\right)  ,
\end{equation}
which, together with eqn. $\left(  \ref{eq:EP}\right)  $, gives the
constraints on the energy of the particle%
\begin{equation}
E_{\text{low}}\equiv\frac{qQ}{r_{+}\left(  M,Q\right)  }<E<M_{e}\left(
Q+q\right)  -M\equiv E_{\text{up}}. \label{eq:AdSECon}%
\end{equation}
Since $q\ll Q$, we can expand $M_{e}\left(  Q+q\right)  $ and obtain
\begin{equation}
E_{\text{up}}\simeq M_{e}\left(  Q\right)  -M+M_{e}^{\prime}\left(  Q\right)
q+\frac{M_{e}^{\prime\prime}\left(  Q\right)  }{2}q^{2},
\end{equation}
where
\begin{equation}
M_{e}^{\prime}\left(  Q\right)  =\frac{Q}{r_{e}\left(  Q\right)  }\text{ and
}M_{e}^{\prime\prime}\left(  Q\right)  =\frac{\sqrt{\sqrt{1+12Q^{2}/l^{2}}-1}%
}{l\sqrt{2/3+8Q^{2}/l^{2}}}>0.
\end{equation}
If the initial black hole is extremal, the lower and upper bounds on $E$
become%
\begin{equation}
E_{\text{low}}=\frac{qQ}{r_{e}\left(  Q\right)  }\text{ and }E_{\text{up}%
}\simeq M_{e}^{\prime}\left(  Q\right)  q+\frac{M_{e}^{\prime\prime}\left(
Q\right)  }{2}q^{2}>E_{\text{low}}.
\end{equation}
So there always exists a test changed particle with $E_{\text{low}%
}<E<E_{\text{up}}$, which can overcharge the extremal RN-AdS black hole. For a
near-extremal RN-AdS black hole with $Q$ and $M=M_{e}\left(  Q\right)
+\epsilon^{2}$, the lower and upper bounds on $E$ become%
\begin{equation}
E_{\text{low}}\simeq M_{e}^{\prime}\left(  Q\right)  q-A\left(  Q\right)
q\epsilon\text{ and }E_{\text{up}}\simeq M_{e}^{\prime}\left(  Q\right)
q+\frac{M_{e}^{\prime\prime}\left(  Q\right)  }{2}q^{2}-\epsilon^{2},
\end{equation}
where $A\left(  Q\right)  >0$ is some function of $Q$, and $\epsilon$ is a
small parameter. To have $E_{\text{up}}>E_{\text{low}}$, we find
\begin{equation}
q\equiv a\epsilon>\frac{-A\left(  Q\right)  +\sqrt{A^{2}\left(  Q\right)
+2M_{e}^{\prime\prime}\left(  Q\right)  }}{M_{e}^{\prime\prime}\left(
Q\right)  }\epsilon. \label{eq:AdSNEQ}%
\end{equation}
The constraints $\left(  \ref{eq:AdSECon}\right)  $ gives the energy $E$ of
the particle,%
\begin{equation}
E=M_{e}^{\prime}\left(  Q\right)  a\epsilon+b\epsilon^{2}\text{ with
}-A\left(  Q\right)  a<b<\frac{M_{e}^{\prime\prime}\left(  Q\right)  }{2}%
a^{2}-1. \label{eq:AdSNEE}%
\end{equation}
Therefore, a charged particle with its charge and energy satisfying eqns.
$\left(  \ref{eq:AdSNEQ}\right)  $ and $\left(  \ref{eq:AdSNEE}\right)  $,
respectively, can overcharge the near-extremal black hole. In summary, WCCC is
always violated for extremal and near-extremal RN-AdS black holes.

\section{RN Black Hole in a Cavity}

\label{Sec:RNBHC}

In this section, we throw a charged particle into a RN black hole enclosed in
a cavity and test the first and second laws of thermodynamics and WCCC. We now
consider a thermodynamic system with a RN black hole enclosed in a cavity, the
wall of which is at $r=r_{B}$. The $4$-dimensional RN black hole solution is%
\begin{align}
ds^{2}  &  =-f\left(  r;M,Q\right)  dt^{2}+\frac{dr^{2}}{f\left(
r;M,Q\right)  }+r^{2}\left(  d\theta^{2}+\sin^{2}\theta d\phi^{2}\right)
\text{,}\nonumber\\
f\left(  r;M,Q\right)   &  =1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}\text{, }%
A=A_{t}\left(  r\right)  dt=-\frac{Q}{r}dt\text{,}%
\end{align}
where $M$ and $Q$ are the black hole charge and mass, respectively. The
Hawking temperature $T_{\text{BH}}$ of the black hole is given by%
\begin{equation}
T_{\text{BH}}=\frac{1}{4\pi}\frac{\partial f\left(  r;M,Q\right)  }{\partial
r}|_{r=r_{+}\left(  M,Q\right)  }=\frac{1}{4\pi r_{+}\left(  M,Q\right)
}\left(  1-\frac{Q^{2}}{r_{+}^{2}\left(  M,Q\right)  }\right)  ,
\label{eq:CavityTBH}%
\end{equation}
where $r_{+}\left(  M,Q\right)  =M+\sqrt{M^{2}-Q^{2}}$ is the radius of the
outer event horizon. Suppose that the wall of the cavity is maintained at a
temperature of $T$. It was shown in \cite{IN-Braden:1990hw} that the system
temperature $T$ can be related to the black hole temperature $T_{\text{BH}}$
as%
\begin{equation}
T=\frac{T_{\text{BH}}}{\sqrt{f\left(  r_{B};M,Q\right)  }}\text{,}
\label{eq:CavityT}%
\end{equation}
which means that $T,$ measured at $r=r_{B}$, is blueshifted from
$T_{\text{BH}}$, measured at $r=\infty$. The thermal energy $\mathcal{E}$ and
potential $\Phi$ of this system were \cite{IN-Braden:1990hw}%
\begin{align}
\mathcal{E}  &  =r_{B}\left[  1-\sqrt{f\left(  r_{B};M,Q\right)  }\right]
,\nonumber\\
\Phi &  =\frac{A_{t}\left(  r_{B}\right)  -A_{t}\left(  r_{+}\right)  }%
{\sqrt{f\left(  r_{B};M,Q\right)  }}. \label{eq:CavityEphi}%
\end{align}
The physical space of $r_{+}\left(  M,Q\right)  $ is bounded by%
\begin{equation}
r_{e}\left(  Q\right)  \leq r_{+}\left(  M,Q\right)  \leq r_{B}\text{,}
\label{eq:rBound}%
\end{equation}
where $r_{e}\left(  Q\right)  =Q$ is the horizon radius of the extremal black hole.

After we throw a particle of energy $E$ and charge $q$ into the RN black hole,
the thermal energy and charge of the system are changed from $\left(
\mathcal{E}\text{, }Q\right)  $ to $\left(  \mathcal{E+}d\mathcal{E}\text{,
}Q+dQ\right)  $. The energy and charge conservation give
\begin{align}
dQ  &  =q\text{,}\nonumber\\
d\mathcal{E}  &  =\frac{1}{\sqrt{f\left(  r_{B};M,Q\right)  }}\left(
dM-\frac{QdQ}{r_{B}}\right)  =E, \label{eq:CavitydQdE}%
\end{align}
where we use eqn. $\left(  \ref{eq:CavityEphi}\right)  $ to express
$d\mathcal{E}$ in terms of $dM$ and $dQ$. Here we assume that the radius
$r_{B}$ of the cavity is fixed during the absorption. Eqn. $\left(
\ref{eq:CavitydQdE}\right)  $ leads to the variation of the black hole mass
$M$,
\begin{equation}
dM=\sqrt{f\left(  r_{B};M,Q\right)  }E+\frac{qQ}{r_{B}}. \label{eq:CavitydM}%
\end{equation}
To discuss the thermodynamic laws, the final RN black hole after absorbing the
particle is assumed to possess an event horizon, which is located at
$r=r_{+}\left(  M,Q\right)  +dr_{+}$. Similar to the RN-AdS case, we have%
\begin{equation}
\frac{\partial f\left(  r;M,Q\right)  }{\partial r}|_{r=r_{+}\left(
M,Q\right)  }dr_{+}+\frac{\partial f\left(  r;M,Q\right)  }{\partial
M}|_{r=r_{+}\left(  M,Q\right)  }dM+\frac{\partial f\left(  r;M,Q\right)
}{\partial Q}|_{r=r_{+}\left(  M,Q\right)  }dQ=0\text{.}%
\end{equation}
Eqns. $\left(  \ref{eq:EP}\right)  $, $\left(  \ref{eq:CavityTBH}\right)  $,
$\left(  \ref{eq:CavityT}\right)  $ and $\left(  \ref{eq:CavitydM}\right)  $
then give
\begin{equation}
TdS=\left\vert P^{r}\left(  r_{+}\left(  M,Q\right)  \right)  \right\vert
+\frac{qQ}{r_{+}\left(  M,Q\right)  }-q\Phi, \label{eq:CavityTdS}%
\end{equation}
where $S=\pi r_{+}^{2}\left(  M,Q\right)  $ is the entropy of the system.
Using eqn. $\left(  \ref{eq:CavityEphi}\right)  $, we rewrite eqn. $\left(
\ref{eq:CavityTdS}\right)  $ as
\begin{equation}
TdS=\left\vert P^{r}\left(  r_{+}\left(  M,Q\right)  \right)  \right\vert
+\left(  \frac{1}{r_{+}}-\frac{1}{r_{+}\sqrt{f\left(  r_{B};M,Q\right)  }%
}+\frac{1}{r_{B}\sqrt{f\left(  r_{B};M,Q\right)  }}\right)  QdQ,
\label{eq:CavityTdSL}%
\end{equation}
where the prefactor of $QdQ$ is positive. It can show that, when $T=0$ (i.e.,
$Q=M$), both sides of $\left(  \ref{eq:CavityTdSL}\right)  $ are zero, which
cannot provide any information about $dS.$ For a non-extremal black hole, eqn.
$\left(  \ref{eq:CavityTdSL}\right)  $ gives that the entropy increases when
$dQ>0$. However when $dQ<0$, the entropy can increases or decrease depending
the value of $dQ$. So the second law of thermodynamics is indefinite for a RN
black hole in a cavity. Substituting eqn. $\left(  \ref{eq:EP}\right)  $ into
eqn. $\left(  \ref{eq:CavityTdS}\right)  $, we obtain%
\begin{equation}
d\mathcal{E}=\Phi dQ+TdS\text{,}%
\end{equation}
which is the first law of thermodynamics.

To overcharge a RN black hole of mass $M$ and charge $Q$ in a cavity by a test
particle of energy $E$ and charge $q$, the mass $M+dM$ and charge $Q+q$ of the
final configuration must satisfy%
\begin{equation}
M+dM<Q+q, \label{eq:CavityMQ}%
\end{equation}
which, due to eqn. $\left(  \ref{eq:CavitydM}\right)  $, puts an upper bound
on $E$,%
\begin{equation}
E<E_{\text{up}}\equiv\frac{Q+q-M-\frac{qQ}{r_{B}}}{\sqrt{f\left(
r_{B};M,Q\right)  }}.
\end{equation}
On the other hand, eqn. $\left(  \ref{eq:EP}\right)  $ also puts a lower bound
on $E$,%
\begin{equation}
E>\frac{qQ}{r_{+}\left(  M,Q\right)  }\equiv E_{\text{low}}\text{.}%
\end{equation}
For an extremal black hole with $M=Q$, we find%
\begin{equation}
E_{\text{up}}=E_{\text{low}}=q\text{,}%
\end{equation}
which indicates that the extremal black hole cannot be overcharged.
Considering a near-extremal RN black hole with $Q$ and $M=Q+\epsilon^{2}$, one
can show that%
\begin{equation}
E_{\text{low}}\simeq q-\frac{\sqrt{2}q\epsilon}{\sqrt{Q}}\text{ and
}E_{\text{up}}\simeq q-\frac{\epsilon^{2}}{1-Q/r_{B}}.
\end{equation}
If $E_{\text{up}}>E_{\text{low}}$, the charge $q$ of the test particle should
be bounded from below,%
\begin{equation}
q\equiv a\epsilon>\frac{\epsilon}{\left(  1-Q/r_{B}\right)  }\sqrt{\frac{Q}%
{2}}. \label{eq:Cavityq}%
\end{equation}
Moreover, the corresponding energy $E$ of the particle is
\begin{equation}
E=a\epsilon+b\epsilon^{2}\text{ with }-\frac{\sqrt{2}a}{\sqrt{Q}}<b<-\frac
{1}{1-Q/r_{B}}. \label{eq:CavityE}%
\end{equation}
So a test particle with $\left(  q,E\right)  $ in the parameter region
$\left(  \ref{eq:Cavityq}\right)  $ and $\left(  \ref{eq:CavityE}\right)  $
can overcharge the near-extremal RN black hole in a cavity, which invalidates
the WCCC.

\section{Conclusion}

\label{Sec:Con}

In this paper, via absorbing a test charged particle, we calculated the
variations of thermodynamic quantities of RN black holes with two boundary
conditions, namely the asymptotically AdS boundary and the Dirichlet boundary
in the asymptotically flat spacetime. With these variations, we checked the
first and second laws of thermodynamics and WCCC in these two cases. Our
results are summarized in Table \ref{tab:1}. In the limit of $l\rightarrow
\infty$, a RN-AdS black hole becomes a RN black hole. Similarly, a RN black
hole in a cavity also reduces to a RN black hole when $r_{B}\rightarrow\infty
$. We find that taking the limits of our results in both AdS and cavity cases
gives the same result about the validity of the thermodynamic laws and WCCC
for a RN black hole, which is also presented in Table \ref{tab:1}.

\begin{table}[tbh]
\centering%
\begin{tabular}
[c]{|p{0.6in}|p{1.6in}|p{1.95in}|p{1.95in}|}\hline
& RN-AdS BH & RN BH in a cavity & RN BH\\\hline
1st Law & Satisfied. & Satisfied. & Satisfied.\\\hline
2nd Law & Satisfied. & Satisfied for $dQ>0$. Indefinite for $dQ<0$. &
Satisfied.\\\hline
WCCC & Violated for extremal and near-extremal BHs. & Satisfied for extremal
BH. Violated for near-extremal BH. & Satisfied for extremal BH. Violated for
near-extremal BH.\\\hline
\end{tabular}
$\ \ $\caption{{\small Results for the first and second laws of thermodynamics
and weak cosmic censorship conjectures (WCCC), which are tested for a RN-AdS
black hole, a RN black hole in a cavity and a RN black hole by absorbing a
test charged particle.}}%
\label{tab:1}%
\end{table}

In \cite{IN-Carlip:2003ne,IN-Lundgren:2006kt}, thermodynamic and phase
structure of a RN-AdS black hole and a RN black hole in a cavity were shown to
be strikingly similar. However, it was found that thermodynamic geometry in
these two cases behaves rather differently \cite{IN-Wang:2019cax}, which
implies that there may be a connection between the black hole microstates and
the boundary condition. In this paper, we showed that the validity of the
second law of thermodynamic and WCCC in the AdS and cavity cases are quite
different, which further motivates us to explore the connection between the
internal microstructure of black holes and the boundary condition.

To understand the scale of energy required for a particle to overcharge a
black hole, we can convert physical quantities in Plack units to SI units. In
fact, for a particle of charge $q$ and energy $E$ in Plack units, the charge
and the energy in SI units are $qq_{p}$ and $Em_{p}c^{2}$, respectively, where
$q_{p}\equiv\sqrt{4\pi\epsilon_{0}\hbar c}=e/\sqrt{\alpha}$ is the Planck
charge, $\epsilon_{0}$ is the permittivity of free space, $e$ is the
elementary charge, $\alpha$ is the fine structure constant, and $m_{p}%
=\sqrt{\hbar c/G}$ is the Planck mass. From eqns. $\left(  \ref{eq:AdSNEQ}%
\right)  $, $\left(  \ref{eq:AdSNEE}\right)  $, $\left(  \ref{eq:Cavityq}%
\right)  $ and $\left(  \ref{eq:CavityE}\right)  $, we find that a test
particle that can overcharge a charged black hole should have
\begin{equation}
E\sim q\text{ (in Planck units),}%
\end{equation}
which leads to%
\begin{equation}
E\sim\frac{\left\vert q\right\vert }{e}\left(  \sqrt{\alpha}m_{p}c^{2}\right)
\sim\frac{\left\vert q\right\vert }{e}\times10^{18}%
%TCIMACRO{\unit{GeV}}%
%BeginExpansion
\operatorname{GeV}%
%EndExpansion
\text{ (in SI units).}%
\end{equation}
For example, if one throws an ionized Hydrogen nuclei of $m\sim1%
%TCIMACRO{\unit{GeV}}%
%BeginExpansion
\operatorname{GeV}%
%EndExpansion
/c^{2}$ and $q=e$ to overcharge a charged black hole, the nuclei is
ultrarelativistic with an enormous kinetic energy $\sim10^{18}%
%TCIMACRO{\unit{GeV}}%
%BeginExpansion
\operatorname{GeV}%
%EndExpansion
$ to overcome the electrostatic repulsion between the particle and the black hole.

In this paper, we discussed the WCCC in the test limit, in that the
interaction between test particles and the black hole background is neglected.
Although this method is simple and straightforward, it is subject to several
limitations. For example, our results showed that the final mass of the black
hole needs to have a second-order correction in $q$ to overcharge the black
hole. However, if one calculates the mass consistently to order $q^{2}$, the
test limit, which is valid only to linear order in $q$, is not enough, and
hence all second-order effects, e.g., self-force and finite size effects,
should be considered. In \cite{IN-Sorce:2017dst}, a general formula for the
full second-order correction to mass was proposed, and it was found that the
WCCC is valid for Kerr-Newman black holes up to the second-order perturbation
of the matter fields. In this new version of the gedanken experiment, the WCCC
was tested and found to be valid for various black holes
\cite{IN-An:2017phb,IN-Ge:2017vun,Con-Jiang:2019ige,Con-Jiang:2019vww}. In
particular, the WCCC was investigated for RN-AdS black holes in the extended
phase space in \cite{Con-Wang:2019bml}, which showed that the WCCC cannot be
violated for RN-AdS black holes under the second-order approximation of the
matter field perturbations. On the other hand, we used the test limit to study
the WCCC for RN-AdS black holes in the normal phase space with fixed
cosmological constant, and found that the WCCC is violated. Apart from
different phase spaces considered, differences between our results and these
in \cite{Con-Wang:2019bml} suggest that corrections beyond the test limit can
play an important role in the analysis of the WCCC for charged AdS black holes
and black holes in a cavity. We leave this for future work.

\bigskip

\begin{acknowledgments}
We are grateful to Haitang Yang and Zhipeng Zhang for useful discussions and
valuable comments. This work is supported in part by NSFC (Grant No. 11875196,
11375121 and 11005016).
\end{acknowledgments}

\end{document}